# Copying style, Extracting value: Illustrators’ Perception of AI Style Transfer and its Impact on Creative Labor

### v1

@@ 5.2.3. Aesthetic choices are semantically meaningful
For her, the aesthetic manipulation of the image, “distorting” is the way to “say something”. This entanglement of aesthetics and semantics is therefore crucial to illustrators’ style. Our findings indicate that illustrators rely on both semantic and aesthetic elements to convey their style in their work, making the goal of disentangling content from aesthetics useless at best and damaging at worst. As we will show in the discussion, while this conception of style transfer does not benefit artists, it does benefit other stakeholders of the creative industry.
Figure 5. Detailed side by side of participant’s work (Pn) and style transfer results (Pn’)
@@ 5.3. Style as an emergent quality of creative work
In this quote, the illustrator outlines what is inherently missing from style transfer, generative models in general, and taste. Comparing his predicament with the changes in typeface availability since the 1980s, he addresses pre-emptively arguments suggesting that quantity will beget quality (larger dataset, bigger compute, etc.) to argue that without taste, style is useless. By centering taste as the X factor of illustration work, he gives the key to the shortcomings raised so far in style transfer. Taste ties texture, colors, and other low-level aesthetic elements to their context and ties the semantic aspect of an image with the aesthetic choices made to depict it. By performing a version of style that relies on its crystallization, extraction, and application regardless of context, generative models preclude the possibility for artists to use them in meaningful ways.
@@
Figure 6. Images generated with the control prompt ”the last supper”
[…]
@@ 5.4. Differential perceptions of the output
[P1]: “It’s really interesting to think about how this tool is reshaping what we think about or making us confront how we see ourselves, how other people see us and what different people pick up, or attribute value to”
@@
Crucially, we found that the perception of the model’s success was aligned with the participant’s proximity to the style evaluated, i.e., how familiar they were to it. While, as we have shown, none of the participants felt that the model was really successful in copying their own style, they were also more likely to find other participants’ output more successful:
[P1 about P3]: “This looks like yeah, it’s… they got him”
@@
Conversely, the other participant, looking at the same images, found the results surprisingly successful, down to the exact aspects of mark-making and brush strokes the first participants deemed missing:
@@
6. Discussion
Connecting their perception of style transfer’s results and limitations to the challenges they already face in their industry (outlined in section 2.2), we discuss our findings in the context of the larger political economy of creative labor that constitutes illustrators’ reality.
First, we show how the logic of style transfers matches clients’ desire to extract value from illustrations in cheaper and faster ways. We then examine how style transfer can be seen as a supply-chain optimization tool rather than a creativity support tool. We conclude with limitations and implications for future research.
@@ 6.1. The political economy of style transfer
As our findings suggest, style transfer generates a sort of “shallow copy” of style, which different actors can see differently as successful or not. Here, we ask who in the illustration industry might this shallow copy benefit?
6.1.1. Copying style, extracting value
For clients, styles are resources that can be extracted from one context, crystalized in a moodboard, then reapplied elsewhere without conserving any traces of its source. Style transfer fits neatly in this context, and participants often made digressions about their professional work during our interviews. Some artists have likened the model’s performance of style to the already existing appropriation of artists’ work in the industry:
@@ 6.1.1. Copying style, extracting value
This “grab” of illustration (a striking term echoing the larger ”data grab” of tech companies (Mejias and Couldry, 2024)) by clients as we are discussing style transfer models points to how both practices share a similar logic. This match is evidenced by how some clients, constantly seeking to optimize efficiency and lower costs, increasingly use generative AI as a blueprint for how to work with illustrators:
@@
6.1.2. Image analogies, image homologies
Our findings also points to how AI systems lack the structural nuances and processes that allow her to adapt how her style is implemented in response to a client’s brief and instead lead with superficial aesthetics. This echoes a nuance that Aaron Hertzmann introduced in his seminal work on “image analogies” (Hertzmann et al., 2001). In biology, analogous elements share functional similarities but no structural kinship (like the wings of a bat and those of a bee). Homologous elements, on the other hand, might look different but share structural similarities that point to common ancestors (the human hand and the bat wing, for example) (Wilkins et al., 2014). What style transfer generates is, for some, analogous to artists’ style, but only on the surface. Unfortunately, this nuance between analogies and homologies has been lost in discussions of style transfer since Hertzmann, which led to hyperbolic statements that take artists’ styles and models’ results as equivalent.
Yet, for all the reasons outlined in our findings, the way illustrators create style is not comparable to what the models can generate. For them, style is about process and structure, which aesthetics follow. Continuing the previous quote, this illustrator explained:
@@ 6.1.2. Image analogies, image homologies
Since the release of generative AI models, many professional artists have reported a loss of income, with clients eager to switch to cheaper options (of Authors, 2024). While this displacement and replacement of creative labor by AI solutions seemingly points to an equivalence between the two, our study shows this is not true. What these models generate in terms of style is not similar to the result or process of artists’ work. Style transfer systems, we suggest, participate actively in such displacement by offering a useful tool not to artists but to clients.
@@ 6.2. From creativity support to supply chain optimization tool
From web scrapping (Lavigne, 2023), and the constitution of datasets (Denton et al., 2021; Crawford and Paglen, 2021; Scheuerman et al., 2021) to the bias introduced by data workers (Miceli and Posada, 2022; Sap et al., 2022) or the output (Wevers, 2018), it has been shown how machine learning and politics go hand in hand. As Langdon Winner (Winner, 2017) pointed out, the politics of technology do not only reside in its use but also in its very design. While many style transfer papers mention the potential misuse of their system, we suggest that the design of a style transfer task inherently fits the objectifying strategies of corporations seeking to benefit from artistic labor, regardless of its use.
Returning to our understanding of style as a boundary object, we suggest that the version of style performed by style transfer, while preventing artists from finding it useful, is a very productive boundary object for the capitalist extractivist strategies of clients. Thus, style transfer aligns with the corporate interests of a managerial class (e.g., marketing teams) eager to derive value from artistic work without working with or remunerating artists. In that sense, extracting style from artists participates in broader processes of extracting value from labor.
@@
Artists have always been an unruly class of workers under capitalism, their work unwieldy for the predictable, risk-averse needs of capitalist production (Bedard, 2020; Child, 2019; Beech, 2015). As a result, the creative industries have designed several strategies (moodboards, in-house copyists or work-for-hire contracts) to facilitate the extraction of value from their work. Similarly, style transfer systems are designed with the assumption that style is a neutral, discreet object that can be extracted from one source and applied to another. Together, these strategies and technologies constitute a socio-technical system designed to optimize value extraction from artistic work, which “metastasizes capitalism” (Keyes et al., 2019). Our analysis of style transfer contributes to approaches calling for “located accountabilities” of system design (Suchman, 2002). Designing systems involves creating and managing boundary objects that facilitate or hinder collaboration between various communities and can enhance power imbalance. Researching style cannot be limited to what Donna Haraway (Haraway, 1988) has called “the god trick of seeing everything from nowhere” but has to be grounded in a critical understanding of the “normative grounds” (Li et al., 2023) that generative models create for their users.
@@
Our study shows that looking at (and working with) style demands a standpoint and that designers should be aware of how, as a boundary object, it “resists any single, totalizing, or universal point of view” (Bardzell, 2010). Furthermore, taking the perspectives of commercial artists like illustrators has helped compensate for the cultural blindspot of studying art in isolation from other social systems like economy or politics (Child, 2019). If, as anthropologist Lily Chumley (Chumley, 2016) writes, “style is one way to know you are in capitalism,” style transfer is, under the garb of a creativity support tool, essentially a form of supply-chain optimization.
6.3. Limitations and Implications for Future Work
@@
Our study is limited to the perspectives of a small sample of professional illustrators working in editorial, branding and advertising contexts. While this choice was essential to spend enough time with each artist to explain the project, answer concerns, and conduct long interviews and follow-ups, a larger sample would provide more nuanced results. We introduced the concept of “boundary objects” to study style, which opens doors for more research on the perceptions of other actors working with style in HCI. For the scope of this study, we centered illustrators’ perspectives and discussed other actors’ relation to style indirectly through their eyes. This was important to foreground illustrators’ narratives in a technology that impacts them most, but our theoretical frameworks would gain from a greater variety of views. Finally, we limited our study to one form of style transfer, the fine-tuning of a StableDiffusion model with a LoRa. Other systems might provide additional insights into the performance and perception of different style transfers. We want to stress that our research did not focus on the performance of a particular model, and the participants’ evaluation of style transfer pertained more to its underlying logic (e.g., Content-Style disentanglement) common to most style transfer tasks.
@@
These limitations are invitations for future research to study style in all its complexity and pluralism. While we limited our sample to illustrators in the editorial, advertising, and branding industries, illustrators working in publishing, entertainment, or animation would undoubtedly enrich these perspectives. Additionally, we invite future research on style transfer to integrate the economic value of style in their study to overcome the limitations of the cultural divide between art and commerce, which does not reflect reality and limits usability.
@@ 7. Conclusion
The concept of style is incredibly valuable to illustrators, both artistically and economically. Interestingly, their perception of style in generative AI research often goes unexamined, even when it is central to the systems’ design. This oversight has led some designers of generative AI models to assert, hyperbolically, that models can replicate “artistic styles”. While style transfer results might be analogous to style (i.e. they share superficial similarities with it) (Hertzmann et al., 2001), we’ve shown they are not homologous to it (i.e. they lack its structural substance). However, the conflation between the two has created a misleading equivalence between machines’ and artists’ creations, which has been used to justify illustrators’ displacement from the creative industry [^3] See for example OpenAI’s CTO Mira Murati’s infamous statement about generative AI replacing creative jobs that ”shouldn’t have been there in the first place.”: https://www.youtube.com/watch?v=Ru76kAEmVfU.
@@
How can we explain, then, that generative AI models are nonetheless disrupting the creative industry? By understanding style transfer’s results as multifaceted boundary objects, we showed that while useless to artists, this shallow copy of style might be enough for their clients. For the artists we interviewed, this is nothing new. The ability of generative AI models to extract value from their work while not crediting or remunerating them merely follows a long history of similar practices in their industry that predate AI. Generative AI, in general, and style transfer, in particular, have often been praised as innovative technological solutions, but for illustrators who try to make a living from their art, they mostly carry on old capitalist problems.

### v2

@@ 5.2.3. Aesthetic choices are semantically meaningful
For her, the aesthetic manipulation of the image, “distorting” is the way to “say something”. This entanglement of aesthetics and semantics is therefore crucial to illustrators’ style. Our findings indicate that illustrators rely on both semantic and aesthetic elements to convey their style in their work, making the goal of disentangling content from aesthetics useless at best and damaging at worst. As we will show in the discussion, while this conception of style transfer does not benefit artists, it does benefit other stakeholders of the illustration industry.
Figure 7. Detailed side-by-side of participant’s work (Pn) and style transfer results (Pn’)
@@
pairs of 2 images are presented on a grid. Each pair is composed of 2 illustrations. The first image of each pair is created by one of the participants on the left-hand side, and four illustrations are generated by our model on the right-hand side. Above the illustrator’s four images is their participant code (Pn), and above the generated images is Pn’. Each pair gives an overview of general stylistic similarity between the illustrators’ work and their models.
@@ 5.3. Style as an emergent quality of creative work
In this quote, the illustrator outlines what is inherently missing from style transfer, generative models in general, is taste. The topic of taste and generative AI exceeds the scope of this paper (for a survey of the topic, see (Hullman et al., 2023)), but we address it here as a more general form of aesthetic evaluation. Comparing his predicament with the changes in typeface availability since the 1980s, the illustrator addresses pre-emptively arguments suggesting that quantity will beget quality (larger dataset, bigger compute, etc.) to argue that without taste, style is useless. By centering taste as a key element of illustration work, he gives the key to the shortcomings raised so far in style transfer. Taste ties texture, colors, and other low-level aesthetic elements to their context and ties the semantic aspect of an image with the aesthetic choices made to depict it. By performing a version of style that relies on its crystallization, extraction, and application regardless of context, generative models preclude the possibility for artists to use them in meaningful ways.
@@ 5.4. Differential perceptions of the output
[P1]: “It’s really interesting to think about how this tool is reshaping what we think about or making us confront how we see ourselves, how other people see us and what different people pick up, or attribute value to.”
@@
Crucially, we found that the perception of the model’s success was aligned with the participant’s proximity to the style evaluated, i.e., how familiar they were to it. While, as we have shown, none of the participants felt that the model was really successful in copying their own style, they were also more likely to find other participants’ output to look like their work:
[P1 about P3]: “This looks like yeah, it’s… they got him.”
@@
Conversely, the other participant, looking at the same images, found the results surprisingly accurate, down to the exact aspects of mark-making and brush strokes the first participant deemed missing:
@@ 5.5. The political economy of style transfer
As our findings suggest so far, style transfer generates a sort of “shallow copy” of style, which different actors can see differently as successful or not. We now turn to how illustrators linked what was generated by the model to how their clients relate to style.
Participants often made digressions about their professional work during our interviews and some artists have likened the model’s performance of style to the already existing appropriation of artists’ work in the industry:
@@
This “grab” of illustration (a striking term echoing the larger “data grab” of tech companies (Mejias and Couldry, 2024)) by clients as we are discussing style transfer models points to how both practices share a similar logic. This match is evidenced by how some clients, constantly seeking to optimize efficiency and lower costs, increasingly use generative AI as a blueprint for how to work with illustrators:
@@
Yet, for all the reasons outlined in our findings, the way illustrators create style is not comparable to what the models can generate. For them, style is about process and structure, which aesthetics follow. While clients might be tempted to conflate an artist’s work with an AI’s output, the same illustrator explained:
@@
Since the release of generative AI models, many professional artists have reported a loss of income, with clients eager to switch to cheaper options (of Authors, 2024). While this displacement and replacement of stylistic labor by AI solutions seemingly points to an equivalence between the two, our study shows this is not true. What these models generate in terms of style is not similar to either the result or process of artists’ work.
6. Discussion
On the one hand, style transfer research suggests that AI models are able to generate images in the style of illustrators, implying that there is some sort of equivalence between the two. On the other hand, our findings show that illustrators, when given the opportunity to use these models, are unimpressed by its output. How can we make sense of this tension? In this section, we discuss this paradox by analyzing the social context of style transfer, and particularly its political economy.
Connecting their perception of style transfer’s results and limitations to the challenges they already face in their industry (outlined in section 2.2), we discuss our findings in the context of the larger political economy of illustration as a profession, which constitutes illustrators’ reality. First, we introduce the concept of “boundary objects” to understand how style, rather than a single “thing” to be extracted from the data, is a complex multifaceted object seen and mobilized differently by various communities. We then examine how style transfer creates a partial analogy to artists’ style. Finally we show style transfer can be seen as a supply-chain optimization tool rather than merely a creativity support tool.
6.1. Style as boundary object
Style sits at the intersection of many stakeholders (artists, their clients, managers, agents, etc.) who desire to do different things with images. Their different, sometimes competing, goals call for an analysis that can accommodate a multiplicity of views of the same object. To do this, we propose conceptualizing style as a “boundary object.”
Understanding style as a boundary object enables us to address the plurality of style perceptions to question the validity of universalizing metrics to evaluate stylistic similarity. This means considering the designers and users of style transfer as partially interacting with style from different perspectives (Velt et al., 2020; John et al., 2004; Jennings, 2005). Engaging with the critical literature that has questioned the universality of datasets (Crawford and Paglen, 2021; Scheuerman et al., 2021; Denton et al., 2021), output (Wevers, 2018), or the positionality of data workers (Miceli and Posada, 2022; Sap et al., 2022), we use boundary objects to show that what tasks like style transfer generate also embed a standpoint (Bardzell, 2010) [^2] While our aim here is to tend to the social plurality of style, we acknowledge there is an intuitive common perception of it grounded in cognition(Augustin et al., 2008; Cupchik et al., 2009)..
Boundary objects also raise our awareness of power dynamics in the generation, perception, and control of styles. By facilitating or hindering collaboration between different communities, they can contribute to the power imbalance between them (Chung and Adar, 2023). In the case of the style generated by style transfer, we will show that it hinders collaboration with artists but facilitates the capitalist extraction of value from their work by their clients. In doing so, we link style transfer to the political economy of illustration work, responding to feminist and STS scholars’ call to connect technological solutions to their political context and impact (Winner, 2017; Suchman, 2002; Rogaway, 2015).
The semiotics of style are complex and worth a note to clarify what we talk about when we talk about style as a boundary object. Style can refer both to an abstract idea, a “generative principle” (Wilf, 2013) that guides the production of artworks; and the concrete occurrences, the artworks that embody this idea in the visible world. In Charles S. Peirce (Peirce, 1974) terminology, these levels are known as the distinction between two levels of signs, type and token. Types are ideal and abstract, for example, the idea of the letter “A.” Tokens, on the other hand, are the concrete occurrences we encounter in the world, e.g., the lowercase “a,” the capitalized “A,” or a hand-drawn “a.” Thus various tokens can point to the same type. Both are, in practice, inseparable, as the type can only exist, socially, through its tokens (Wetzel, 2018), e.g., we only get a sense of “a style” through the perception of images conveying it [^3] For an analysis of style and computation in semiotic terms, see (Wilf, 2013).. For our analysis, this entanglement between abstract types and concrete tokens is important as it is exactly because of this that style can function as a boundary object. It can function both as an abstract common object to various communities but in its concrete, localized actualization, it differs widely, creating tensions and synergies.
6.2. A partial version of style
As Langdon Winner (Winner, 2017) pointed out, the politics of technology do not only reside in its use but also in its very design. While many style transfer papers mention the potential misuse of their system, we suggest that the design of a style transfer task inherently fits the objectifying strategies of corporations seeking to benefit from artistic labor, regardless of its use.
As evidenced by the findings in section 5.4, the evaluation of the success of style transfer is socially situated, with ingroup differences of perception noticeable even in such a small sample as ours. Moving away from approaches that would dismiss such differences as self-serving biases, we suggest that these findings points to how style, as a boundary object, can be seen differently by various actors, while still being understood as common to all. In this study, illustrators have not found the model to be useful for a variety of reasons: style transfer only generates limited similarities with artists’ work (section 5.1), is limited to the disentanglement of style and content (section 5.2), and creates a fixed snapshot of aesthetics rather than enables the emergence of style within creative practice (section 5.3). These limitations point to how, rather than a single object that is either style or not, AI-generated styles are but one version of what style can be. It is only one facet of a larger, more complex object.
Note that this nuanced understanding of computation styles was present in the early days of style transfer. For example, this echoes a distinction that Aaron Hertzmann introduced in his seminal work on “image analogies” (Hertzmann et al., 2001). In biology, analogous elements share functional similarities but no structural kinship (like the wings of a bat and those of a bee). Homologous elements, on the other hand, might look different but share structural similarities that point to common ancestors (the human hand and the bat wing, for example) (Wilkins et al., 2014). What style transfer generates is, for some stakeholders, in some contexts, analogous to artists’ style, but only on the surface. Unfortunately, this careful nuance between analogies and homologies has since been lost in discussions of style transfer, which led to hyperbolic statements that artists’ styles and models’ results are equivalent.
To understand how such a partial version of style can seemingly have such a disruptive impact on the illustration industry, we must move away from style as a mere aesthetic problem and return to the political economy of illustration, and who profits from the extraction of style from images, if not artists.
@@ 6.3. From creativity support to supply chain optimization tool
Artists have always been an unruly class of workers under capitalism, their work unwieldy for the predictable, risk-averse needs of capitalist production (Bedard, 2020; Child, 2019; Beech, 2015). As a result, the creative industries have designed several strategies (moodboards, in-house copyists or work-for-hire contracts) to facilitate the extraction of value from their work. Similarly, style transfer systems are designed with the assumption that style is a neutral, discreet object that can be extracted from one source and applied to another. Together, these strategies and technologies constitute a socio-technical system designed to optimize value extraction from artistic work, which “metastasizes capitalism” (Keyes et al., 2019).
Under the same term “style”, AI-generated styles now exist alongside human-created ones, a shared social space increasingly resulting in competition between the two. This is the “robust” quality of style as a boundary object that allows it to be common to several communities. However, as we have seen, with regards to artists, AI-generated style is actually a boundary-closing object, it hinders collaboration with them, generating output at odds with their work both in terms of process and results. For clients on the other hand, the object becomes boundary-opening, it is aligned with their goals and expectations of what images are for and how to extract value from them. When left unquestioned as a single, unified object, style becomes the ground on which artists’ obsolescence is discussed, another form of human labor that can simply be automated. But when analyzed as a multifaceted boundary object, what appeared to be neutral becomes particular, design choices become political, aesthetic results become economic. It is thus every researcher’s responsibility to be attuned to the boundary they open and close when designing systems. Our analysis of style transfer contributes to approaches calling for “located accountabilities” of system design (Suchman, 2002). Designing systems involves creating and managing boundary objects that facilitate or hinder collaboration between various communities and can enhance power imbalance. Researching style cannot be limited to what Donna Haraway (Haraway, 1988) has called “the god trick of seeing everything from nowhere” but has to be grounded in a critical understanding of the “normative grounds” (Li et al., 2023) that generative models create for their users.
@@
Our study shows that looking at (and working with) style demands a standpoint and that designers should be aware of how, as a boundary object, it “resists any single, totalizing, or universal point of view” (Bardzell, 2010). Using common sense terms such as “style” does not mean every stakeholder is understanding the same thing, which in turn means that the systems we design do not necessarily work for everyone. While the model’s output failed as a boundary object for creative work for artists, they suggested it might succeed as a boundary object for their clients to extract value from their work, which is two very different things. Taking the perspectives of commercial artists like illustrators has helped compensate for the cultural blindspot of studying art in isolation from other social systems like economy or politics (Child, 2019). By analyzing style transfer from the vantage point of the political economy of illustration, we have shown that style, rather than simply about aesthetics, is akin to a form of labor from which value can be extracted. We thus relocate discussions of style transfer alongside other works on the exploitation of the “unwitting labourer” (Morreale et al., 2023) and micro-workers (Tubaro et al., 2020) required to feed machine learning models. In a context where so many Creativity Support Tools might displace workers from their industry, we encourage future HCI research on creative AI to engage directly with questions of labor. If, as anthropologist Lily Chumley (Chumley, 2016) writes, “style is one way to know you are in capitalism,” style transfer is, under the garb of a creativity support tool, essentially a form of supply-chain optimization.
7. Limitations and Implications for Future Work
@@
Our study is limited to the perspectives of a small sample of professional illustrators working in editorial, branding and advertising contexts who agreed to have their work fine-tuning a model. We acknowledge that this strong self-selection bias limits the insights and reproducibility of our results. While this choice was essential to spend enough time with each artist to explain the project, answer concerns, and conduct long interviews and follow-ups, a larger sample would provide more nuanced results. We introduced the concept of “boundary objects” to study style, which opens doors for more research on the perceptions of other actors working with style in HCI. For the scope of this study, we centered illustrators’ perspectives and discussed other actors’ relation to style indirectly through their eyes. This was important to foreground illustrators’ narratives in a technology that impacts them most, but our theoretical frameworks would gain from a greater variety of views. Finally, we limited our study to one form of style transfer, the fine-tuning of a StableDiffusion model with a LoRa. Other systems might provide additional insights into the performance and perception of different style transfers. We want to stress that our research did not focus on the performance of a particular model, and the participants’ evaluation of style transfer pertained more to its underlying logic (e.g., Content-Style disentanglement) common to most style transfer tasks.
@@
These limitations are invitations for future research to study style in all its complexity and pluralism. While we limited our sample to illustrators in the editorial, advertising, and branding industries, illustrators working in publishing, entertainment, or animation would undoubtedly enrich these perspectives.
Our study opens several directions for future research. We suggest that because art is not separate from other social spheres, especially economy (Menger, 2014; Becker, 2023), and the creative industries at large rely on individuals commodifying some personal traits (style, voice, appearance, etc.), research in creative AI should always engage with how systems benefit from, align with or disrupt the local power dynamics in a given field. How are other artistic fields (music, acting, etc.) organizing this interface between the personal practice of artists and the value they generate for their industry? How does our study echo more direct forms of AI-mediated extraction of even more personal objects such as actors’ bodies or singers’ voices? How does a commitment to understanding such dynamics change the way and for whom systems are designed? Conversely, what kinds of socio-technical structure enable artists to use AI tools in their work in their own terms (Caramiaux and Fdili Alaoui, 2022)?
As we alluded to, style transfer optimizes a view of illustration aligned with the largely noncreative (at least in the romantic sense) managerial class of the creative industry, rather than artists’. Future research at the intersection of creative AI and HCI could therefore focus on this class of workers to study how they shape the supply chain of style as a commodity. How do they interpret style, its source, role, and value? What makes the difference between a system aligned with artists and one aligned with managers? Future research may also want to study artists’ uptake of these systems within their own practice, but also in relationship of the political economy of such practice. Understanding style as a boundary object opens such directions by localizing it within larger social contexts, and offers HCI researchers a more concrete approach to an otherwise abstract concept.
@@ 8. Conclusion
The concept of style is incredibly valuable to illustrators, both artistically and economically. Interestingly, their perception of style in generative AI research often goes unexamined, even when it is central to the systems’ design. This oversight has led some designers of generative AI models to assert, hyperbolically, that models can replicate “artistic styles”. While style transfer results might be analogous to style (i.e., they share superficial similarities with it) (Hertzmann et al., 2001), we’ve shown they are not homologous to it (i.e., they lack its structural substance). However, the conflation between the two has created a misleading equivalence between machines’ and artists’ creations, which has been used to justify illustrators’ displacement from the creative industries [^4] See for example OpenAI’s CTO Mira Murati’s infamous statement about generative AI replacing creative jobs that “shouldn’t have been there in the first place.”: https://www.youtube.com/watch?v=Ru76kAEmVfU.
@@
How can we explain, then, that generative AI models are nonetheless disrupting the creative industries? By understanding style transfer’s results as multifaceted boundary objects, we showed that while useless to artists, this shallow copy of style might be enough for their clients. For the artists we interviewed, this is nothing new. The ability of generative AI models to extract value from their work while not crediting or remunerating them merely follows a long history of similar practices in their industry that predate AI. Generative AI, in general, and style transfer, in particular, have often been praised as innovative technological solutions, but for illustrators who try to make a living from their art, they mostly carry on old capitalist problems.